• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jin

    Log entry time 21:06:12 on November 11, 2008

    Entry number 248236

    This entry is a followup to: 248222

    keyword=Target Holding field messed up by wired Bigbite setting, Pol -> 10%, target -> empty, test pumping up with new laser

    It appears someone turned Bigbite magnet ON around 3:30PM without Halog or letting target people know. Even worse, the Bigbite current is set to 100A, rather than standard 710A, which caused a false field correction.

     

    Bigbite field change caused target field shift by 3G, which basically kicked NMR peak to the edge of sweep, and introduced high polarization loss. Pol =~10% now!

     

    Target is moved back to empty location to start pumping. Bigbite field is reset to 0A.

     

    See also,

    Today’s NMR sweep report

    A previous normal NMR



    A copy of this log entry has been emailed to: jpchen@jlab.org, camsonne@jlab.org, yqiang@jlab.org